OpenStudy (anonymous):

Consider the circle of radius 5 centered at (0,0). Find an equation of the line tangent to the circle at the point (3,4) in slope intercept form.

OpenStudy (bradely):

Slope , m = (4-0)/(3-0) =4/3 Perpendicular slope =-3/4 Equation of tangent y=(-3/4)x+c point(3,4) 4=(-3/4)(3)+c c=4+(9/4) =25/4 y=(-3/4)x+25/4 source: http://www.mathskey.com/question2answer/